OpenStudy (anonymous):

Need help: Find the exact value of sec(2tan^-1 (3/4)

OpenStudy (jdoe0001):

hmmm lemme take a poke

OpenStudy (jdoe0001):

$$\bf sec\left(2tan^{-1}\left(\cfrac{3}{4}\right)\right)\\ sec(2\theta) \implies \cfrac{1}{cos(2\theta)}\\ cos(2\theta) = 1- 2 sin^2(\theta)\\ \text{now let's check above}\\ \color{green}{ tan^{-1}\left(\cfrac{3}{4}\right) \implies \cfrac{b}{a}\\ c = \sqrt{4^2+3^2} = 5 } $$

OpenStudy (jdoe0001):

so \(\bf cos(2\theta) = 1- 2 sin^2(\theta) \implies cos(2\theta) = 1- 2\left(\cfrac{3}{5}\right)^2\)

OpenStudy (jdoe0001):

so, that'd give you the cosine, flip it around to the secant

OpenStudy (jdoe0001):

gives me 7/25, so the secant will be 25/7